OpenStudy (anonymous):

just making sure of my answer :) log[4]16(32)^1/2 is the answer x=9/4?

Parth (parthkohli):

\[\log_4{16} + \log_4{32^{1/2}} = 2 + \frac{1}{2}\left(\log_416 + \log_4 2\right) = 2 + \frac{1}{2}\left(2 + \frac{1}{2}\right) = 2 + 1 + \frac{1}{4} = \frac{13}{4}\]